A local police station receives an average 6.3 emergency telephone calls per hour. These calls are Poisson distributed.

The probability that the station will get at least 3 calls per hour is:

a.         0.0397

  1.       0.0941
  2.       0.9059
  3.       0.9502
  4.       0.9630
Answer #1

Let say X~Poisson(\lambda)

So,

PMF

P(X=x) = e^{- \lambda } \frac{\lambda ^x}{x!} , x=0,1,2,.....

Thus Mean of X:

E(X)= \sum_{x=0}^{\infty}xP(X=x) \\\\ = \sum_{x=0}^{\infty}xe^{- \lambda } \frac{\lambda ^x}{x!} \\\\ = e^{- \lambda } \lambda \sum_{x=1}^{\infty}\frac{\lambda^{x-1}}{(x-1)!}.........at x=0 the term is 0

= e^{- \lambda } \lambda \sum_{y=0}^{\infty}\frac{\lambda^{y}}{(y)!}.....................y=x-1

= e^{- \lambda } \lambda e^{\lambda }..................by the formula of taylor series

=\lambda

Thus the mean or average =\lambda

For this problem ,

\lambda =6.3

X: Number of local calls per hour

X~Poisson(6.3)

So, P(at least 3 calls per hour) = P(X>=3) :

P(X\geq 3) =1- P(X<3) \\\\ =1- \sum_{x=0}^{2} e^{-6.3} \frac{6.3^x}{x!} \\\\ = 1- 0.04984 =0.95016 \approx 0.9502

Option "d"